# Rethinking Detection Based Table Structure Recognition for Visually Rich Documents

### 01 Dec 2023 v1

@@ 4.4.1 Metrics misalignment
4.4.2 Comparison with Sparse R-CNN
Besides the metrics misalignment, our other important observation is that the Deformable-DETR and Sparse R-CNN also achieved promising results even though they are not as good as the proposed method. Therefore, using Sparse R-CNN as an example, we further analyze the difference between our proposed TSRDet and Sparse R-CNN and discuss how Sparse R-CNN implicitly deals with the multi-label definition.
Sparse R-CNN mainly consists of a backbone network, learnable region proposal, and a series of Dynamic Heads containing Dynamic Instance Interactive Head. Therefore, for a input image $x$ with initialized region proposal feature embedding $e_{pro}$ and corresponding proposal boxes $b$, the inputs to the $head_{cls}$ and $head_{reg}$ in the first Dynamic Heads can be generated by Equations 910,  11 and 12, in which $BACKBONE,ROI\_POOL,SELF\_ATTN$ and $PROJ$ are the backbone network, ROI_Pooling, self attention and projection layers. Since Sparse R-CNN also utilizes a cascade architecture, the $N$th Dynamic Head takes the outputs of $(N-1)$th Head’s regression result and $e_{dyn}$ as the proposal boxes and proposal feature embedding. Since Sparse R-CNN has multiple ways to initialize the region proposal boxes, such as Random boxes, Image Size boxes, and Grid boxes, we follow the setting of Section 4 to assume the region proposal boxes are initialized with Image Size boxes, which means that for an input image $x$, its proposal boxes $b$ can be treated as constants. Therefore, the classification and regression results are determined by the learned region proposal feature embedding $e_{pro}$, which can differ even when their corresponding boxes are identical. Therefore, Sparse R-CNN can naturally deal with the multi-label definition for the TSR task.
$e_{roi}=ROI\_POOL(BACKBONE(x),b)$ (9)
$e_{dyn}=DYN\_CONV(SELF\_ATTN(e_{pro},e_{pro},,e_{pro}),e_{roi})$ (10)
$e_{cls}=PROJ_{cls}(e_{dyn})$ (11)
$e_{reg}=PROJ_{reg}(e_{dyn})$ (12)
By contrast, for the classification head of Cascade R-CNN, their inputs are features extracted by applying RoI Pooling operation, as defined by Equation 2. In other words, the input features of the classification head are determined by the bounding boxes once the model is trained. Therefore, there can only be one classification result for a proposal box, making Cascade R-CNN challenging to deal with the multi-label definition for the TSR task.
@@ 5 Conclusion and Future Work
In this study, we propose TSRDet, a tailored Cascade R-CNN based solution for the TSR task, directly detecting six types of table components. The experimental results show that the proposed solution outperforms different types of state-of-the-art models, including image-to-sequence and graph-based models. Considering the definition of six components, a bounding box can simultaneously be a Row, Projected Row Header, and Column Header, which means it is a multi-label classification. We refactor the multi-label classification to regular single-label classification by generating pseudo-classes. Besides, the proposed Spatial Attention module can be inserted into the backbone network to build long-range dependencies, so deformable convolution is applied to further improve the model performance. Since the proposed method is detection-based and focuses on well-formatted, visually rich documents, one major limitation is that it may fail to deal with irregular tables, such as rotated and distorted tables. Integrating instance segmentation with detection models is included in our future research agenda to deal with irregular tables because instance segmentation can handle the irregular shape and be guided by the bounding boxes.

### 10 Jan 2024 v2

@@ 5.4.2 The Misalignment of Metrics
5.4.3 Deformable Convolution and Spatial Attention
As discussed in sections 3.4 and 3.5, both generating good local features and building long-range dependencies are essential for a detection-based TSR model, and deformable convolution can improve the local feature generation but has the risk leading to the over-optimization to the detection performance. The ablation study’s experimental results can somewhat demonstrate our analysis. Considering the performance of Ablation 1 with deformable convolution in Tables 8 and 9, its TEDS is 84.35%, lower than the Cascade R-CNN baseline (87.49%), but its mAP is improved from 95.23% to 97.22%. These results not only show the misalignment of COCO and TEDS metrics but also demonstrate that merely improving local features can make the model fit empty spaces better, as shown in Figures 9(a), but does not help alleviate the multi-label detection issue. Therefore, deformable convolution needs to be applied with other methods. On the other hand, our proposed Spatial Attention Module can improve the mAP and structure-only TEDS simultaneously if we compare the performance of Ablation 3 and 4, and also can be used with deformable convolution together to improve the structure-only TEDS further, as shown in Figure 8, demonstrating the effectiveness of building long-range dependencies.
5.4.4 Other Observations
Besides the observations discussed in previous sections, the experimental results also show other phenomena that can be helpful in our model design. One observation is that Cascade R-CNN has better detection performance on small objects than Sparse R-CNN. For example, on the FinTabNet dataset, $APs$ of Sparse R-CNN is only 62.11%, while the Cascade R-CNN baseline and our proposed TSRDet reach 87.32% and 91.60%. This phenomenon might be caused by their methods of generating regional proposals. As discussed in section 3.1, Cascade R-CNN uses RPN to generate regional proposals, which regress and classify anchor boxes, and the anchor boxes are generated by sliding the pre-defined boxes with different aspect ratios and sizes on the feature map of multiple scales. Therefore, Cascade R-CNN uses a dense proposal generation method sun2021sparse with more region proposals, meaning that Cascade R-CNN can use the parameters of RPN to generate more high-quality small region proposals. By contrast, Sparse R-CNN uses sparse learnable regional proposals to replace dense proposals generated by the RPN, which can avoid parameter tuning of RPN but limit its performance on small objects. Another interesting observation is that the baseline Cascade R-CNN can work better on complex tables than simple tables, which is very different from other benchmark models. This phenomenon is caused by the fact that the spanning cells in complex tables are usually in the Column Row Headers, which can alleviate the multi-label detection issue. For example, Figures 2(c) and Figures 2(d) show two samples from PubTables1M dataset, in which the former is a complex table and the latter is a simple table. Because of the existence of Spanning Cells in Figure 2(c), the Column Header does not share its bounding box with any rows, which avoids multi-label detection. By contrast, the sample in Figure 2(d) does not contain any Spanning Cell, making its Column Header share its bounding box with a Row, which is the challenging multi-label detection to Cascade R-CNN. As comparisons, Deforamble-DETR and Sparse R-CNN can deal with multi-label detection, and their performance on simple tables is better than complex tables regarding the structure-only TEDS, as shown in Tables 34and 5.
@@ 6 Conclusion and Future Work
In this study, we first revisit existing detection-based TSR solutions and analyze the critical design aspects for a successful detection-based TSR model, including the problem formulation, the characteristics of detection models, and the characteristics of TSR tasks. Our analysis can be a guideline for improving the performance of a detection-based model. To demonstrate our analysis and findings, we propose TSRDet by applying simple methods to tailor the Cascade R-CNN, which can outperform different types of state-of-the-art models, including image-to-sequence and graph-based models. Even though we only applied very simple methods to a tow-stage detection model, there should be other methods to further improve the model based on our analysis. For example, vision transformers can be considered to build long-range dependencies. Transformer-based detection models, such as Sparse R-CNN, can also be considered as base models with the benefits of dealing with multi-label detection tasks and learnable proposals. Besides, since the proposed method is detection-based and focuses on well-formatted, visually rich documents, one major limitation is that it may fail to deal with irregular tables, such as rotated and distorted tables. Integrating instance segmentation with detection models can be another direction to deal with irregular tables because instance segmentation can handle irregular shapes and be guided by bounding boxes.